Two forces of 200 N each are applied to an object,
Answer #1

8)
Fnet of 2 given force = 200*cos 70 + 200*cos 70
= 137 N
so 3rd force should be equal in magnitude of Fnet but opposite in direction
F3 = 137 N
Answer: a

9)
mass of earth, M = 5.972*10^24 Kg
Radius of earth, R = 6371000 m
r = (6371000 + 230000) = 6601000 m
use:
G*M*m/r^2 = m*v^2/r
G*M/r = v^2
v^2 = (6.673*10^-11)*(5.972*10^24) / (6601000 )
v = 7769.9 m/s
T = 2*pI*r/v
= 2*pi*6601000 / 7769.9
= 5337.9 s
= 89 min
Answer: d

I am allowed to answer only 1 question at a time but I have answered 2 questions here